OpenStudy (anonymous):

PLEASE HELP: Determine the correct equation for the following verbal sentence: The ticket price, t, for a concert attended by 220 people equals the total revenue, r, divided by the number of people who attended.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

t = r/220

OpenStudy (anonymous):

Translate the sentence into an equation using n to represent the unknown number. Then solve the equation for n. a number subtracted from 46 is -21 @kelliegirl33

OpenStudy (anonymous):

46 - n = 21 46 - 21 = n 25 = n

OpenStudy (anonymous):

Write the following phrase as an algebraic expression. Use n for “a number.” the product of two and the sum of a number and eight @kelliegirl33

OpenStudy (anonymous):

2(n + 8)

OpenStudy (anonymous):

The distance an object travels is determined by how fast it goes and for how long it moves. If an object’s speed remains constant, the formula is d=r*10, where d is the distance, r is the speed, and t is the time. If you drive at 50 miles per hour for 5 hours, how far will you have traveled? a. 200 miles b. 10 miles c. 250 miles d. 300 miles @kelliegirl33

OpenStudy (anonymous):

d = r * t (this is the distance- rate- time formula r = 50 t = 5 you can do this, can't you ?

OpenStudy (anonymous):

250

OpenStudy (anonymous):

you got it :)